# YOLO-LLTS: Real-Time Low-Light Traffic Sign Detection via Prior-Guided Enhancement and Multi-Branch Feature Interaction

### v2

@@ IV-C Comparisons with the State-of-the-arts
Tab. II presents a performance comparison of different traffic sign detection models on the CNTSS dataset. Our model achieves an accuracy of 88.3%, surpassing the second-ranked YOLO-TS by 1.0%. In terms of recall, our approach reaches 74.9%, which is 1.0% higher than the second-ranked GOLD-YOLO-L, demonstrating superior detection coverage. Our model’s F1 score is 81.0, outperforming GOLD-YOLO-L by 1.5. The mAP50 of our model is 81.2%, surpassing GOLD-YOLO-L by 1.3%, and the mAP50-95 is 60.1%, which is 1.9% higher than YOLO-TS. With a parameter count of 13.9M and an FPS of 82.0, these results indicate that our approach maintains high accuracy and robustness across different detection difficulties.
@@ IV-D Ablation Study
TABLE IV: Ablation study for different components on CNTSSS. The best result is marked in blue.
  Model HRFM-TOD PGFE DFEDR mAP50 mAP50:95 FPS
@@
IV-D1 Effectiveness of HRFM-TOD
To validate the effectiveness of HRFM-TOD, we retain only this module on top of the baseline. As shown in the table, the HRFM-TOD module improves the mAP50 from 75.1% to 77.6%, an increase of 2.5%, and the mAP50:95 from 53.3% to 55.5%, an increase of 2.2%. The FPS also increases from 75.4 to 77.0. The experiment demonstrates that the HRFM-TOD module not only enhances detection performance but also accelerates model inference speed.
TABLE V: Ablation study for $\gamma$ and $\delta$ on CNTSSS. The best result is marked in blue.
@@ IV-D2 Effectiveness of PGFE
To validate the effectiveness of PGFE, we conducted corresponding ablation experiments. As shown in the table, the PGFE module improves the mAP50 from 75.1% to 78.3%, an increase of 3.2%, and the mAP50:95 from 53.3% to 55.5%, an increase of 2.2%. When PGFE and HRFM-TOD are combined, the mAP50 increases by 4.4%, and the mAP50:95 increases by 5.8%. The experiments demonstrate that the HRFM-TOD module significantly enhances low-quality images, resulting in better detection performance. Additionally, we conducted experiments with different parameter settings in the formula. As shown in the table, the model performs best when $\gamma$ and $\delta$ are set to 2 and 2.5, respectively.
IV-D3 Effectiveness of DFEDR
Since the DFEDR module can only be used in conjunction with the HRFM-TOD module, we validated the combined effect of both modules. As shown in the second to last row of the table, the mAP50 and mAP50:95 increased by 4.5% and 5.1%, respectively, compared to the baseline. When compared to the HRFM-TOD-only configuration, mAP50 and mAP50:95 increased by 2.0% and 2.9%, respectively. The experiment demonstrates that the DFEDR module effectively integrates multi-receptive field features, improving detection performance.
@@ IV-E Error Analysis and Model Deployment
At the same time, we performed inference speed testing on the mobile edge device, NVIDIA Jetson AGX Orin. The experiments were conducted in an Ubuntu 18.04 operating system environment, using the PyTorch deep learning framework (version 2.1.0), with optimization support provided by Jetpack 5.1. We selected 1,500 test images from the CCTSDB-2021 dataset for the speed testing. Without utilizing TensorRT acceleration, the inference time for YOLO-LLTS per image was 44.9 ms, with an FPS of 22.3. The experiments demonstrate that our model exhibits good real-time performance on edge devices. This indicates that our model has promising applications in advanced driver assistance systems (ADAS) and autonomous driving systems. In addition, we evaluated the model using a Mobile Edge Computing Device in real-world road scenarios. In the scene depicted in Fig. 12, traffic signs were successfully detected and accurately classified. This demonstrates the full effectiveness of our model in real-world applications.
@@ V Conclusion
In this paper, we propose YOLO-LLTS, an end-to-end real-time traffic sign detection algorithm specifically designed for low-light environments. To address the lack of nighttime-specific traffic sign datasets, we constructed a novel dataset named Chinese Nighttime Traffic Sign Sample Set (CNTSSS). This dataset includes images captured under varying low-light conditions from dusk to midnight, covering diverse scenarios such as urban, highway, and rural environments under different weather conditions. We introduce the High-Resolution Feature Map for Small Object Detection (HRFM-TOD) module to effectively address the indistinct features of small objects under low-light conditions, significantly improving both detection accuracy and inference speed. Moreover, we design the Multi-branch Feature Interaction Attention (MFIA) module, enabling deep interaction and fusion of features across multiple receptive fields, thereby enhancing the model’s ability to capture and utilize critical information. Furthermore, we propose the Prior-Guided Enhancement Module (PGFE) to alleviate challenges such as increased noise, reduced contrast, and blurriness under low-light environments, significantly boosting the detection performance. Experimental results demonstrate that our approach achieves state-of-the-art performance on the TT100K-night, CNTSSS, and CCTSDB2021 datasets. The deployment experiments on edge devices further validate the practical effectiveness and real-time applicability of our method. In future work, we plan to expand our CNTSSS dataset with more diverse scenarios and further optimize the algorithm to enhance robustness and generalizability for real-world autonomous driving applications. We would also like to release our source code and dataset to facilitate further research in this area.
@@ Acknowledgments
This project is jointly supported by the National Natural Science Foundation of China (Nos.52172350, W2421069, 51775565 and 61003143), the Guangdong Basic and Applied Research Foundation (No. 2022B1515120072), the Guangzhou Science and Technology Plan Project (No.2024B01W0079), the Nansha Key RD Program (No.2022ZD014), the Science and Technology Planning Project of Guangdong Province (No.2023B1212060029).

### v3

@@ IV-C Comparisons with the State-of-the-arts
Tab. II presents a performance comparison of different traffic sign detection models on the CNTSSS dataset. Our model achieves an accuracy of 88.3%, surpassing the second-ranked YOLO-TS by 1.0%. In terms of recall, our approach reaches 74.9%, which is 1.0% higher than the second-ranked GOLD-YOLO-L, demonstrating superior detection coverage. Our model’s F1 score is 81.0, outperforming GOLD-YOLO-L by 1.5. The mAP50 of our model is 81.2%, surpassing GOLD-YOLO-L by 1.3%, and the mAP50-95 is 60.1%, which is 1.9% higher than YOLO-TS. With a parameter count of 13.9M and an FPS of 82.0, these results indicate that our approach maintains high accuracy and robustness across different detection difficulties.
@@
Tab. IV presents a performance comparison of different traffic sign detection models on the GTSDB-night dataset. Our model achieves an accuracy of 77.1%, surpassing the second-ranked YOLO-TS by 4.7%. In terms of recall, YOLO-LLTS reaches 62.4%, which is 3.2% higher than YOLO-TS, demonstrating superior detection coverage. Our model’s F1 score is 69.0, outperforming YOLO-TS by 3.9. Our method also excels in the mAP50 and mAP50-95 metrics, achieving 74.8% and 65.8%, surpassing YOLO-TS by 7.5% and 9.8%. With a parameter count of 9.9M and an FPS of 101.0, these results indicate that our approach maintains high accuracy and robustness across different detection difficulties. The excellent performance on the GTSDB-night dataset demonstrates the applicability of our model to different regions or countries with varying traffic sign styles.
@@ IV-D Ablation Study
TABLE V: Ablation study for different components on CNTSSS. The best result is marked in blue.
[…]
@@
TABLE VI: Ablation study for $\gamma$ and $\delta$ on CNTSSS. The best result is marked in blue.
@@
IV-D1 Effectiveness of HRFM-SOD
To validate the effectiveness of HRFM-SOD, we retain only this module on top of the baseline. As shown in the table, the HRFM-SOD module improves the mAP50 from 75.1% to 77.6%, an increase of 2.5%, and the mAP50:95 from 53.3% to 55.5%, an increase of 2.2%. The FPS also increases from 75.4 to 77.0. The experiment demonstrates that the HRFM-SOD module not only enhances detection performance but also accelerates model inference speed.
@@ IV-D2 Effectiveness of PGFE
To validate the effectiveness of PGFE, we conducted corresponding ablation experiments. As shown in the table, the PGFE module improves the mAP50 from 75.1% to 78.3%, an increase of 3.2%, and the mAP50:95 from 53.3% to 55.5%, an increase of 2.2%. When PGFE and HRFM-SOD are combined, the mAP50 increases by 4.4%, and the mAP50:95 increases by 5.8%. The experiments demonstrate that the HRFM-SOD module significantly enhances low-quality images, resulting in better detection performance. Additionally, we conducted experiments with different parameter settings in the formula. As shown in the table, the model performs best when $\gamma$ and $\delta$ are set to 2 and 2.5, respectively.
IV-D3 Effectiveness of MFIA
Since the MFIA module can only be used in conjunction with the HRFM-SOD module, we validated the combined effect of both modules. As shown in the second to last row of the table, the mAP50 and mAP50:95 increased by 4.5% and 5.1%, respectively, compared to the baseline. When compared to the HRFM-SOD-only configuration, mAP50 and mAP50:95 increased by 2.0% and 2.9%, respectively. The experiment demonstrates that the MFIA module effectively integrates multi-receptive field features, improving detection performance.
@@ IV-D3 Effectiveness of MFIA
Figure 15: Mobile Edge Computing Device NVIDIA Jetson AGX Xavier.
Figure 16: Mobile Edge Computing Device NVIDIA Jetson AGX Orin.
TABLE VII: Testing the detection accuracy and inference times on Jetson AGX Xavier; The first best results are indicated in blue.
  Method Input Size Precision(%) Recall(%) F1 mAP50(%) mAP50:95(%) Param(M)$\downarrow$ Inference time (ms)$\downarrow$ Device YOLOv8-L[64] 640$\times$640 84.4 74.4 79.1 80.5 53.0 43.6 145.8 AGX Xavier YOLOv9-C [65] 640$\times$640 84.8 76.0 80.2 81.9 53.7 50.7 234.0 AGX Xavier YOLOv10-L[66] 640$\times$640 83.9 73.2 78.2 78.7 51.4 25.7 134.2 AGX Xavier YOLOv11-L[67] 640$\times$640 85.8 74.2 79.6 81.2 54.1 25.3 110.9 AGX Xavier YOLOv12-L[68] 640$\times$640 84.9 74.2 79.2 82.8 55.7 26.3 134.8 AGX Xavier Average - 84.2 74.4 79.0 80.4 52.8 34.3 151.9 - YOLO-LLTS(ours) 640$\times$640 88.8$\uparrow$ 81.1$\uparrow$ 84.8$\uparrow$ 87.8$\uparrow$ 57.5$\uparrow$ 10.2 138.8 AGX Xavier
TABLE VIII: Testing the detection accuracy and inference times on Jetson AGX Orin; The first best results are indicated in blue.
  Method Input Size Precision(%) Recall(%) F1 mAP50(%) mAP50:95(%) Param(M)$\downarrow$ Inference time (ms)$\downarrow$ Device YOLOv8-L[64] 640$\times$640 84.4 74.4 79.1 80.5 53.0 43.6 24.6 AGX Orin YOLOv9-C [65] 640$\times$640 84.4 76.0 80.0 81.9 53.7 50.7 55.9 AGX Orin YOLOv10-L[66] 640$\times$640 83.9 73.2 78.2 78.7 51.4 25.7 36.8 AGX Orin YOLOv11-L[67] 640$\times$640 85.8 74.2 79.6 81.2 54.1 25.3 34.7 AGX Orin YOLOv12-L[68] 640$\times$640 84.9 74.2 79.2 82.8 55.7 26.3 48.4 AGX Orin Average - 84.1 74.4 79.0 80.4 52.8 34.3 27.0 - YOLO-LLTS(ours) 640$\times$640 88.8$\uparrow$ 81.1$\uparrow$ 84.8$\uparrow$ 87.8$\uparrow$ 57.5$\uparrow$ 10.2 44.9 AGX Orin
@@ IV-E Error Analysis and Model Deployment
At the same time, we test the detection accuracy and inference times on mobile edge devices, NVIDIA Jetson AGX Xavier and NVIDIA Jetson AGX Orin. As shown in Fig. 15 and Fig. 16, we first connect the device to the display, keyboard, and mouse, then connect the network cable and power supply. We power on the edge computing device and test the trained weights on both the Jetson AGX Xavier and Jetson AGX Orin. The experiments on Jetson AGX Xavier were conducted in an Ubuntu 20.04 operating system environment, with optimization support provided by Jetpack 5.1. The experiments on Jetson AGX Orin were conducted in an Ubuntu 18.04 operating system environment, using the PyTorch deep learning framework (version 2.1.0), with optimization support provided by Jetpack 5.1. We selected 500 nighttime images from the CCTSDB2021 dataset for testing. As shown in Tab. VII and Tab. VIII, YOLO-LLTS achieves the best performance in terms of accuracy. Specifically, our model outperforms the average of YOLOv8 to YOLOv12 by 4.6 in accuracy, 6.7 in recall, and 5.8 in F1 score. It also exceeds the average by 7.4 in mAP50 and 4.7 in mAP50:95. Without utilizing TensorRT acceleration, the number of parameters on the Xavier is 10.2 million, with an inference time of 138.8 ms. On the Orin, the number of parameters is 10.2 million, with an inference time of 44.9 ms. The experiment demonstrates that our model maintains relatively competitive performance even on devices with constrained power.
To evaluate the performance of our model in real-world scenarios, we use the mobile edge device to detect traffic signs on safe roads. We connect the Jetson AGX Orin to the vehicle’s power supply, capture images of traffic signs using the camera, and then conduct testing using the trained weights. In the scene depicted in Fig. 17, traffic signs were successfully detected and accurately classified. This demonstrates the full effectiveness of our model in real-world applications.
@@ V Conclusion
Under low-light conditions, the perception systems of autonomous vehicles are prone to false positives and missed detections, leading to road traffic accidents. In this paper, we propose YOLO-LLTS, an end-to-end real-time traffic sign detection algorithm, which effectively enhances the accuracy and speed of traffic sign recognition, thereby improving the safety of ADAS and autonomous vehicles in low-light environments. To address the lack of nighttime-specific traffic sign datasets, we constructed a novel dataset named Chinese Nighttime Traffic Sign Sample Set (CNTSSS). This dataset includes images captured under varying low-light conditions from dusk to midnight, covering diverse scenarios such as urban, highway, and rural environments under different weather conditions. We introduce the High-Resolution Feature Map for Small Object Detection (HRFM-SOD) module to effectively address the indistinct features of small objects under low-light conditions, significantly improving both detection accuracy and inference speed. Moreover, we design the Multi-branch Feature Interaction Attention (MFIA) module, enabling deep interaction and fusion of features across multiple receptive fields, thereby enhancing the model’s ability to capture and utilize critical information. Furthermore, we propose the Prior-Guided Feature Enhancement Module (PGFE) to alleviate challenges such as increased noise, reduced contrast, and blurriness under low-light environments, significantly boosting the detection performance. Experimental results demonstrate that our approach achieves state-of-the-art performance on the TT100K-night, CNTSSS, CCTSDB2021, and GTSDB-night datasets. The deployment experiments on edge devices further validate the practical effectiveness and real-time applicability of our method.
In future work, we plan to expand our CNTSSS dataset with more diverse scenarios. In addition, although the YOLO-LLTS algorithm shows significant improvements in traffic sign detection and recognition under low-light conditions, it still has limitations in extremely low-light conditions. As shown in Fig. 18, when the road environment has weak lighting or lacks street lamps, our model failed to recognize traffic signs. Future research could further explore ways to enhance the model’s detection performance in extremely low-light or nighttime environments.
Figure 18: A failure case under extreme low-light conditions.
@@ Acknowledgments
This project is jointly supported by the National Natural Science Foundation of China (Nos.52172350, W2421069 and 51775565), the Tongchuang Intelligent Medical Inter-disciplinary Talent Training Fund of Sun Yat-sen University (No.76160-54990001), the Guangdong Basic and Applied Research Foundation (No. 2022B1515120072), the Guangzhou Science and Technology Plan Project (No.2024B01W0079), the Nansha Key RD Program (No.2022ZD014), the Science and Technology Planning Project of Guangdong Province (No.2023B1212060029). We thank Dr. Dezong Zhao from the University of Glasgow and Dr. Kening Li from Shenzhen Institute of Information Technology for their valuable comments.